OpenStudy (anonymous):

Could 2, 4 and 5 be the measures of the sides of a right triangle?

OpenStudy (ipwnbunnies):

Apply the Pythargorean Theorem equation. It only applies to right triangles. \[a^2 + b^2 = c^2\] a and b are the shorter legs of the right triangle. C is the hypotenuse of the right triangle -- which is the longest side and the side opposite the right angle

OpenStudy (neonumbrella5115):

So, using what iPwnBunnies said...\[a^2+b^2=c^2\] \[2^2+4^2=5^2\] Is that equation true?

OpenStudy (anonymous):

No, it isn't. Would that mean that the sides can't make a right triangle?

OpenStudy (anonymous):

Yes.

OpenStudy (neonumbrella5115):

Yup.

OpenStudy (anonymous):

"The square of the hypotenuse has to equal the sum of the squares upon the legs."

OpenStudy (anonymous):

Thank you guys!!